# Few-shot Online Anomaly Detection and Segmentation

### v1

@@ 4.2 Evaluation Protocols
For comprehensively assessing the few-shot learning and online learning capacity of models for dealing with data stream, the evaluation entails both offline and online settings.
@@
Offline evaluation: In the offline setting, the approaches are exclusively trained on the few-shot training dataset and do not utilize data flow for parameter updates. This setting is compared to online evaluating and can assess the effectiveness of the model in learning from online data. Additionally, this evaluation setting can also reflect the ability to learn from few-shot samples to some extent.
@@ 4.3 Evaluation Metrics
0011112222333344445555666677778888999910101010707070708080808090909090100100100100SessionsAccuracy [%](a) CapsuleROCAUC-imageROCAUC-pixelPRO 00111122223333444455556666777788889999101010101111111112121212707070708080808090909090100100100100SessionsAccuracy [%](b) CarpetROCAUC-imageROCAUC-pixelPRO
001111222233334444707070708080808090909090100100100100SessionsAccuracy [%](c) ToothbrushROCAUC-imageROCAUC-pixelPRO 00111122223333444455556666777788889999404040406060606080808080100100100100SessionsAccuracy [%](d) GridROCAUC-imageROCAUC-pixelPRO
00111122223333444455556666777788885050505060606060707070708080808090909090SessionsAccuracy [%](e) TransistorROCAUC-imageROCAUC-pixelPRO 0011112222333344445555666677778888999910101010707070708080808090909090100100100100SessionsAccuracy [%](f) Metal NutROCAUC-imageROCAUC-pixelPRO
@@ 4.4 Comparison results
We present the comparative results of our methods on the MVTec dataset firstly. MVTec is specifically designed for anomaly detection and segmentation, thus we report a complete set of results, including Image-level (ROCAUC %), Pixel-level (ROCAUC %), and per-region overlap (PRO %). To investigate the online learning effects, the variation curve of detection accuracy is depicted in Figure 6.
@@
We then proceed to compare our method with other prevalent approaches in anomaly detection and segmentation. As shown in Table 8, our method demonstrates a significant improvement over previous work, leveraging online learning to achieve an image-level ROCAUC of 87.3 and a PRO score of 87.8. In contrast, approaches lacking online learning capabilities only achieve an image-level ROCAUC of 85.4 and a PRO score of 86.7, as demonstrated by the PatchCore method which utilizes a more intricate neighborhood aggregation feature representation; This outcome underscores the robust online learning capabilities of our approach.
Furthermore, to delve deeper into the performance of online learning, we conducted an evaluation of our method under the offline learning procedure, with the comparison results presented in Table 8. The offline model yields 78.2 and 82.8 in image-level ROCAUC and PRO metrics, though the result is not the best within offline methods, just falling between PaDiM and SPADE, it can achieve a much higher performance after being reinforced by online learning process.
We then modify PaDiM, SPADE, DRAEM and PatchCore to incorporate the capability of online learning: wherein the feature embedding/image is added to the model for parameter updates if the image anomaly score falls below a fixed threshold. PaDiM is almost the same as its offline version, this is because that PaDim divides an image into a grid of $(i,j)\in[1,W]\times[1,H]$ and associates a multivariate Gaussian distribution to fixed position $(i,j)$. But images in MVTec AD dataset are not strictly aligned, which makes the distribution of position $(i,j)$ contains appearance information from different parts of object. This inherent bias results that it is difficult for the model to accurately adjust the distribution parameters through online learning. The online version of SPADE exhibits only a marginal performance improvement. It is important to note that while SPADE shows some progress, the method requires constant enlargement of its memory bank, potentially exceeding device storage limits. This point also proves the effectiveness of our method in another way. DRAEM demonstrates some performance improvement compared to other methods in the comparison, but it still lags significantly behind our approach. DRAEM is a deep neural network model that tends to overfit current data and lose previously learned information [12, 13], which makes it hard to learn a comprehensive image feature distribution. Benefiting from the sophisticated local-aware patch features that aggregation information from neighborhood, PatchCore performs well on offline settings with just a few training samples. However, it also does not improve detection ability through online learning stage. A possible reason for this phenomenon is that this algorithm is easy to sample misjudged outliers into the coreset during online learning. Outliers typically reside at a considerable distance from the normal coreset in the feature space, are more likely to be incorporated into the coreset due to the minimax facility location coreset selection algorithm which selects the farthest point from the current coreset. As a result, the presence of outliers in the coreset notably impacts the learning process and the performance of memory bank-based algorithms.
@@
76767676787878788080808082828282848484848686868688888888909090909292929294949494ImagePRO92.892.892.892.882.282.282.282.288.988.988.988.984.784.784.784.791.191.191.191.176.076.076.076.093.093.093.093.086.586.586.586.5Accuracy [%]PaDimSPADEPatchCoreOurs
@@
Table 2: Mean inference time per image of our method on MVTec AD. Scores are (Image ROCAUC, PRO metric).
@@ 4.5 Inference Time
Inference time is another dimension which we are interested in. We report the results in Table 2. We measure the model inference time on the CPU (Intel i7-4770 @3.4GHz) with a GPU (NVIDIA TITAN X (Pascal)). All methods leverage GPU operations to accelerate the algorithm whenever possible. The statistics, including the online learning stage, are displayed in Table 2. Although our method updates its parameter to enhance accuracy, it just consumes 0.315 seconds more than the algorithm without online learning ability. Generally, our model can process an image within one second while continuously updating itself.
@@ 4.6 Ablation Study
In this section, we firstly investigate the performance under different hyper-parameter settings. The network comprises three hyper-parameters: the initial cluster epoch $N$, the number of the cluster centers $K$, and the maximum age $age_{max}$ of the edge in the network. The hyper-parameters used for comparison with other methods are highlighted with gray background in the table.
The results of hyper-parameter cluster epoch $N$ are reported in Figure 8; we fixed other hyper-parameters and alter the number of the epochs. It is evident that increasing the epoch number $N$ has limited effects on the final result. The per-pixel ROCAUC and per-region overlap are almost identical, while $N$ varies from 10 to 40. The variation range of image-level ROCAUC is slightly more extensive. Notably, even with only one epoch, the performance is quite high, indicating that the initial algorithm has a relatively fast convergence rate.
@@
Image 83.4 86.9 86.3 87.3 87.2
Pixel 93.2 94.7 95.1 95.1 95.3
PRO 83.1 86.8 87.5 87.8 88.3
We then proceed with ablation studies on the size of the initial cluster $K$. The experiment results are presented in Table 3. Generally, the segmentation performance improves with the enlargement of the initial cluster size. The rate of improvement is particularly rapid when the size is small and becomes slower after the size reaches $42\times 42$. Nevertheless, the trend of the image-level ROCAUC is unusual. It reaches an extreme value at $28\times 28$ size and then declines until reaching another one. We will discuss this phenomenon later in Section 5. To strike a balance between accuracy and computational complexity, $K=56\times 56$ is chosen in this paper.
We also research the significance of the maximum age $Age_{max}$ of the edge in the network. Epoch $N$ and the number of the cluster center $K$ are fixed during the studies. The results, shown in Table 4, include Image-level ROCAUC, Pixel-level ROCAUC, and Per-region overlap under a series of $Age_{max}$. Similar to the number of epoch $N$, the Pixel-level ROCAUC and the Per-region overlap is almost the same no matter how $Age_{max}$ changes. Only image-level ROCAUC varies from 86.8 to 87.3 in the research, and after the parameter exceeds 100, the accuracy stabilizes.
@@
Image 87.0 87.1 87.3 87.2 86.9 86.8 86.8
Pixel 95.1 95.1 95.1 95.1 95.1 95.1 95.1
PRO 87.7 87.7 87.8 87.7 87.7 87.8 87.8
Finally, we conduct ablation studies to investigate the effectiveness of the different types of threshold. In this paper, we apply three types of thresholds: the mean distance between the neural node and its neighbors, the maximum distance from the neural node to its neighbors, and the minimum distance. The results are presented in Table 5, obviously, employing the mean distance between the neural node and its neighbors yields the best performance in both detection and segmentation, while the minimum distance achieves the second-highest performance in image-level detection. Unexpectedly, the model without any threshold gets 86.3 in image-level ROCAUC and outperforms the threshold using the maximum distance, which is proposed in video anomaly detection by previous work [57].
@@
Threshold Type Min Mean Max None
Image 87.2 87.3 86.1 86.3
Pixel 94.7 95.1 95.0 95.0
PRO 87.1 87.8 87.6 87.5
@@
Additionally, we give computational information in Table 7. As shown in the table, our method maintains a good balance between performance improvement and parameter quantity. Although SPADE [9] and PatchCore [19] have smaller initial parameter size, they need to expand memory bank size constantly during online learning stage, which poses a risk of storage overflow.
@@ 4.7 Implementation Details
In our experiments, ResNet18 which is pre-trained on the ImageNet dataset is employed as the feature extractor. All the images are resized to $256\times 256$ and then cropped to $224\times 224$. In order to achieve the multi-scale representation, we use features at the end of the first block, the second block, and the third block from the ResNet18 backbone. Feature maps of different sizes are resized to $224\times 224$ by bilinear interpolation. The features from different blocks with the same position are concatenated together and reduced dimension to 100 by random selection. We selected 3136 cluster centers, and the training epoch is set to 10 in the initialization stage of the model. In the online learning stage, the batch size is set to 10. The hyper-parameter $age_{max}$ is set at 25. After obtaining the pixel-level anomaly score, a Gaussian filter ($\sigma=4$) is employed to smooth the map. We reimplemented SPADE, PaDiM, DRAEM and PatchCore under FOADS setting to demonstrate our online performance. ResNet18 is used as the backbone network and the same feature dimension is applied in SPADE, PaDim and PatchCore for a fair comparison.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@ 5 Discussion and Outlook
As mentioned in the ablation study, although there is a certain change trend when the hyper-parameter $K$ varies, the image-level ROCAUC differs. The main reason for this phenomenon is that some sub-classes in the MVTec AD exhibit an opposite trend with changeable hyper-parameter. For example, as shown in 11, in the ablation study of the cluster size $K$, the “Grid” sub-class shows a significant difference compared to others. The peak of “Grid” appears at $28\times 28$ and then declines rapidly, while others are still increasing. This may be attributed to “Grid” just consisting of simple texture patterns, which does not need too many clusters to represent. In the future, it would be beneficial to make the establishment of the model more flexible, for example, the initial size can self-tune to suit different sub-classes. We can begin with only two nodes and determine whether new data belonging to existing clusters. If not, model should grow the scale of the network for better fitting. This iterative growth process enables the model to adapt to the specific requirements of different sub-classes, resulting in an arbitrary-sized network that is tailored to address their respective peculiarities.
@@
Ours Image 99.8 84.2 86.6 94.9 58.6 97.2 98.7 94.1 81.6 61.4 90.6 89.7 81.1 99.6 91.5 87.3
PRO 96.4 77.3 93.4 95.4 79.9 93.3 98.4 83.7 91.4 86.6 78.1 88.1 67.6 94.3 92.9 87.8
6 Conclusion
7 Appendices
@@ Acknowledgements
This work is funded by National Key Research and Development Project of China under Grant No. 2020AAA0105600 and 2019YFB1312000, National Natural Science Foundation of China under Grant No. 62006183 and 62076195, China Postdoctoral Science Foundation under Grant No. 2020M683489, and the Fundamental Research Funds for the Central Universities under Grant No. xzy012020013. We thank MindSpore for the partial support of this work, which is a new deep learning computing framework [^1] https://www.mindspore.cn/.

### v2

@@ 4.2 Evaluation Protocols
For comprehensively assessing the few-shot learning and online learning capacity of models for dealing with data streams, the evaluation entails both offline and online settings.
@@
Offline evaluation: In the offline setting, the approaches are exclusively trained on the few-shot training dataset and do not utilize data flow for parameter updates. This setting is compared to online evaluation and can assess the effectiveness of the model in learning from online data. Additionally, this evaluation setting can also reflect the ability to learn from few-shot samples to some extent.
@@ 4.3 Evaluation Metrics
0111222333444555666777888999101010707070808080909090100100100SessionsAccuracy [%](a) CapsuleROCAUC-imageROCAUC-pixelPRO 0111222333444555666777888999101010111111121212707070808080909090100100100SessionsAccuracy [%](b) CarpetROCAUC-imageROCAUC-pixelPRO
0111222333444707070808080909090100100100SessionsAccuracy [%](c) ToothbrushROCAUC-imageROCAUC-pixelPRO 0111222333444555666777888999404040606060808080100100100SessionsAccuracy [%](d) GridROCAUC-imageROCAUC-pixelPRO
0111222333444555666777888505050606060707070808080909090SessionsAccuracy [%](e) TransistorROCAUC-imageROCAUC-pixelPRO 0111222333444555666777888999101010707070808080909090100100100SessionsAccuracy [%](f) Metal NutROCAUC-imageROCAUC-pixelPRO
@@ 4.4 Comparison results
We first present the comparative results of our methods on the MVTec dataset. MVTec is specifically designed for anomaly detection and segmentation, thus we report a complete set of results, including Image-level (ROCAUC %), Pixel-level (ROCAUC %), and per-region overlap (PRO %). To investigate the online learning effects, the variation curve of detection accuracy is depicted in Figure 6.
@@
We then proceed to compare our method with other prevalent approaches in anomaly detection and segmentation. As shown in Table 3, our method demonstrates a significant improvement over previous work, leveraging online learning to achieve an image-level ROCAUC of 87.4 and a PRO score of 88.1. In contrast, approaches lacking online learning capabilities only achieve an image-level ROCAUC of 85.4 and a PRO score of 86.7, as demonstrated by the PatchCore method which utilizes a more intricate neighborhood aggregation feature representation; This outcome underscores the robust online learning capabilities of our approach. Detailed results for each category in the MVTec dataset are presented in Table 10.
Table 3: Comparison results on MVTec dataset under few-shot online detection setting, we report average results of all sessions: Image (Image-level (ROCAUC %)), PRO (per-region overlap (PRO %))
Offline Methods
Method DRAEM [7] DiffusionAD [11] RegAD-L [41] SPADE [8] PaDim [9] PatchCore  [10] Ours
Image 82.2 84.4 84.9 77.4 77.4 85.4 78.2
PRO 76.6 78.8 85.2 83.1 82.6 86.7 82.8
Online Methods
Image 86.2 86.8 80.1 80.3 77.7 84.9 87.4
PRO 78.9 80.0 81.3 85.1 83.1 85.8 88.1
Furthermore, to delve deeper into the performance of online learning, we conducted an evaluation of our method under the offline learning procedure, with the comparison results presented in Table 3. The offline model yields 78.2 and 82.8 in image-level ROCAUC and PRO metrics, though the result is not the best within offline methods, just falling between PaDiM and SPADE, it can achieve a much higher performance after being reinforced by the online learning process.
[…]
@@
767676787878808080828282848484868686888888909090929292949494ImagePRO92.892.892.882.282.282.288.988.988.984.784.784.791.191.191.176.076.076.093.093.093.086.586.586.5Accuracy [%]PaDimSPADEPatchCoreOurs
@@
Table 4: Mean inference time per image and GPU memory footprint of our method on MVTec AD. Scores are (Image ROCAUC, PRO metric).
@@ 4.5 Inference Time
Inference time is another dimension which we are interested in. We report the results in Table 4. We measure the model inference time on the CPU (Intel i7-4770 @3.4GHz) with a GPU (NVIDIA TITAN X (Pascal)). All methods leverage GPU operations to accelerate the algorithm whenever possible. The statistics, including the online learning stage, are displayed in Table 4. Although our method updates its parameters to enhance accuracy, it just consumes 0.315 seconds more than the algorithm without online learning ability. Generally, our model can process an image within one second while continuously updating itself. Besieds, shown in the Table, our method only consumes around 2.2GB GPU memory, which can run on most consumer-grade GPU cards.
@@ 4.6 Ablation Study
In this section, we first investigate the performance under different hyper-parameter settings. The network comprises three hyper-parameters: the initial cluster epoch $N$, the number of the cluster centers $K$, and the maximum age $age_{max}$ of the edge in the network. The hyper-parameters used for comparison with other methods are highlighted with gray background in the table.
The results of hyper-parameter cluster epoch $N$ are reported in Figure 8; we fixed other hyper-parameters and altered the number of the epochs. It is evident that increasing the epoch number $N$ has limited effects on the final result. The per-pixel ROCAUC and per-region overlap are almost identical, while $N$ varies from 10 to 40. The variation range of image-level ROCAUC is slightly more extensive. Notably, even with only one epoch, the performance is quite high, indicating that the initial algorithm has a relatively fast convergence rate.
@@
Image 82.6 87.4 86.8 87.4 87.1
Pixel 93.7 95.2 95.3 95.4 95.5
PRO 82.9 86.9 87.9 88.1 88.2
We then proceed with ablation studies on the size of the initial cluster $K$. The experiment results are presented in Table 5. Generally, the segmentation performance improves with the enlargement of the initial cluster size. The rate of improvement is particularly rapid when the size is small and becomes slower after the size reaches $\times 42$ 42. Nevertheless, the trend of the image-level ROCAUC is unusual. It reaches an extreme value at $\times 28$ 28 size and then declines until reaching another one. We will discuss this phenomenon later in Section 5. To strike a balance between accuracy and computational complexity, $K=56\times 56$ is chosen in this paper.
We also research the significance of the maximum age $Age_{max}$ of the edge in the network. Epoch $N$ and the number of the cluster center $K$ are fixed during the studies. The results, shown in Table 6, include Image-level ROCAUC, Pixel-level ROCAUC, and Per-region overlap under a series of $Age_{max}$. Similar to the number of epoch $N$, the Pixel-level ROCAUC and the Per-region overlap are almost the same no matter how $Age_{max}$ changes. Only image-level ROCAUC varies from 86.8 to 87.4 in the research, and after the parameter exceeds 100, the accuracy stabilizes.
@@
Image 86.6 87.1 87.4 87.2 86.9 86.7 86.7
Pixel 95.3 95.3 95.4 95.4 95.4 95.4 95.4
PRO 87.7 87.9 88.1 88.1 88.0 88.0 87.9
We also conduct ablation studies to investigate the effectiveness of the different types of thresholds. In this paper, we apply four types of thresholds: the mean distance between the neural node and its neighbors, the median distance to the neighbors, the maximum distance from the neural node to its neighbors, and the minimum distance. The results are presented in Table 7, obviously, employing the median distance between the neural node and its neighbors yields the best performance in both detection and segmentation, while the mean distance achieves the second-highest performance in image-level detection. Unexpectedly, the model without any threshold gets 86.3 in image-level ROCAUC and outperforms the threshold using the maximum distance, which is proposed in video anomaly detection by previous work [42]. These results demonstrate the superiority of our threshold system.
@@
Threshold Type Min Median Mean Max None
Image 87.2 87.4 87.3 86.1 86.3
Pixel 94.7 95.4 95.1 95.0 95.0
PRO 87.1 88.1 87.8 87.6 87.5
@@
Additionally, we give parameter volume information in Table 9. As shown in the table, our method maintains a good balance between performance improvement and parameter quantity. Although SPADE [8] and PatchCore [10] have smaller initial parameter sizes, they need to expand memory bank size constantly during the online learning stage, which poses a risk of storage overflow.
@@ 4.7 Implementation Details
In our experiments, ResNet18 which is pre-trained on the ImageNet dataset is employed as the feature extractor. All the images are resized to $\times 256$ 256 and then cropped to $\times 224$ 224. In order to achieve the multi-scale representation, we use features at the end of the first block, the second block, and the third block from the ResNet18 backbone. Feature maps of different sizes are resized to $\times 224$ 224 by bilinear interpolation. The features from different blocks with the same position are concatenated together and reduced dimension to 100 by random selection. We selected 3136 cluster centers, and the training epoch is set to 10 in the initialization stage of the model. In the online learning stage, the batch size is set to 10. The hyper-parameter $age_{max}$ is set at 25. After obtaining the pixel-level anomaly score, a Gaussian filter ($\sigma=4$) is employed to smooth the map. We reimplemented SPADE, PaDiM, DRAEM, DiffusionAD and PatchCore under FOADS setting to demonstrate our online performance. ResNet18 is used as the backbone network and the same feature dimension is applied in SPADE, PaDim, RegAD and PatchCore for a fair comparison. And environment with CUDA 10.2, Ubuntu 16.04, Python 3.7 is employed for the experiments in this paper.
@@ 5 Discussion and Outlook
5.1 Failure Cases and Limitations
Shown in Figure 11, when evaluated on the MVTec AD dataset, our model does not perform well on narrow and elongated type defects.This is mainly due to the presence of downsampling in convolutional neural networks. The embeddings in the feature map represent a patch rather than a pixel. Thus, the segmentation results are not accurate enough to detect very minor anomalies.
Our method also has some limitations on the selection of initial cluster size. As mentioned in the ablation study, although there is a certain change trend when the hyper-parameter $K$ varies, the image-level ROCAUC differs. The main reason for this phenomenon is that some sub-classes in the MVTec AD exhibit an opposite trend with changeable hyper-parameter. For example, as shown in Figure 12, in the ablation study of the cluster size $K$, the “Grid” sub-class shows a significant difference compared to others. The peak of “Grid” appears at $\times 28$ 28 and then declines rapidly, while others are still increasing. This may be attributed to “Grid” just consisting of simple texture patterns, which does not need too many clusters to represent.
Figure 11: Visualization of failure cases. Left column: Input Images containing anomaly regions. Center column: Ground truth regions of anomalies. Right column: Anomaly Score maps for each test image predicted by our algorithm.
5.2 Outlook
In the future, it would be beneficial to make the establishment of the model more flexible, for example, the initial size can be self-tuned to suit different sub-classes. We can begin with only two nodes and determine whether new data belongs to existing clusters. If not, the model should grow the scale of the network for better fitting. This iterative growth process enables the model to adapt to the specific requirements of different sub-classes, resulting in an arbitrary-sized network that is tailored to address their respective peculiarities.
@@ 7 Appendices
Ours Image 99.8 87.0 84.4 94.2 59.4 97.8 98.4 96.8 80.9 60.8 88.6 89.9 82.8 99.0 91.2 87.4
PRO 96.5 76.2 93.5 95.5 81.2 92.7 98.4 82.8 90.3 87.3 78.5 87.6 72.9 94.6 92.9 88.1
@@ Acknowledgements
This work was supported by National Key R&D Program of China under Grant No. 2021ZD0110400, the Fundamental Research Funds for the Central Universities No.xxj032023020.and sponsored by the CAAI-MindSpore OpenFund, developed on Openl Community.